OpenStudy (anonymous):

Answer please i'll give a medal. What is the third term of this multiplication pattern? multiplication pattern using the formula 3 • 7n – 1 A. 1 B. 21 C. 147 D. 441

OpenStudy (jdoe0001):

\(\large \begin{array}{llll} term({\color{brown}{ n}})&value \\\hline\\ 1& 3 \cdot 7{\color{brown}{ (1)}} - 1\\ 2&3 \cdot 7{\color{brown}{ (2)}} - 1\\ 3&3 \cdot 7{\color{brown}{ (3)}} - 1 \end{array}\)

OpenStudy (jdoe0001):

though I don't see it to be any of the given choices which means the pattern or the choices are off